HUBBLE SPACE TELESCOPE DAILY REPORT #2195 PERIOD COVERED: 0000Z (UTC) 08/25/98 - 0000Z (UTC) 08/26/98 Daily Status Report as of 238/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed NIC/1/2/3 8060 (Observation of the Unusual BL Lac Object/Compact Symmetric Object PKS 1413+135) The Near Infrared Camera (Cameras One, Two and Three) was used to make high resolution observations of the bright core of BL Lac object PKS1413+135. This object is known to be a strong radio source. The observations were completed as scheduled, and no anomalies were noted. 1.4 Completed Four Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.5 Completed Six Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.6 Completed Two Set of STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.7 Completed NIC/1/2/3 7869 (The Morphology of Dense Gas in Seyferts , Obstruction and Fueling of AGNs) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the Seyfert galaxy NGC-5643 to study the morphology of its dense gas. Near-Infrared color maps and images in H-2 emission will be used to trace hot and cold dense gas in this Seyfert galaxy. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed NIC/2 7860 (The Optical Emission Line Nebulae of High Redshift Radio Galaxies) The Near Infrared Camera (Camera Two) was used to make high resolution observations of the optical nebula in the high redshifted radio galaxy 4C+19.71. This is to compare the sizes and detailed structures of the nebulae to the radio and visual continuum morphologies on sub-galactic size scales, and relate them to those found in lower redshift radio galaxies studied with HST and ground-based observatories over the same wavelength region. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed Three Sets of STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.10 Completed NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the star HD349726. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.11 Completed STIS/CCD/MA1/MA2 7931 (Scattered Light in the Echelle Modes) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to study scattered light in the STIS Echelle Modes by making observations of the O-class star CPD59D260. The effect is strongest at the shortest red wavelengths. A wavelength calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.12 Completed Four Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.13 Completed NIC/2 7833 (Multiplicity Survey of Alpha Persei: Studying the Evolution and Effects of Companions) The Near Infrared Camera (Camera Two) was used to make observations of the star Alpha-Persei. Extensive observations have established that binaries are twice as common amongst pre-main sequence stars than older solar neighborhood G-dwarfs. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed STIS/CCD/MA1 8016 (Effect of MAMA Charge Offsetting on Sensitivity and Dispersion Accuracy - Repeat) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations of G191B2B. In order to conserve charge on the STIS MAMA microchannel plates and maintain even detector quantum efficiency, the STIS MAMA spectroscopic modes are differentially offset each month along and across the detector. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 10 Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 19 Successful : 19 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was reset four times. There was another session of the IP test. This iteration included passes scheduled via TDRS-East and STGT, and ran from 237/2027Z until 237/2053Z. The STOCC processed 32 kbps"Q" channel data, 4 kbps NSSC-1 status buffer dump "I" channel data, and 1.024 Mbps SSR science data playback. FIVEBLK commanding was also performed. No problems were reported. At 237/221956Z, a NICMOS single bit error occurred. NICMOS memory was dumped at 237/2249Z per an open operations request and the upper limit was changed to six. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH